# DataRater: Meta-Learned Dataset Curation

### v1

@@ 5 Conclusion
This approach offers a meaningful step towards automating and improving current dataset curation practices, which currently rely heavily on manual tuning and handcrafted heuristics. Our method provides a scalable and principled way to determine the value of individual data, and shows considerable compute efficiency gains for filtering multiple real-world datasets when training on models of significant scale – thus providing a first proof of concept that meta-learning the data curation pipeline for modern LLMs is possible.
@@ Acknowledgements
The authors would like to express their gratitude to John E. Reid, Paul Michel, Katie Millican, Andrew Dai, Andrei A. Rusu, Jared Lichtarge, Chih-Kuan Yeh, Oriol Vinyals, the Gemini Data Team and the RL team for fruitful discussions, valuable feedback, and support throughout the project.
@@ Appendix A Appendix
Figure 9: Relative performance improvement induced by DataRater models trained on populations consisting of different sizes of inner models, as a function of the meta-training checkpoint. The underlying training dataset is C4/noclean and all evaluations are done with a $50\%$ discard fraction. The relative performance improvement is calculated with respect to a 1B baseline trained on the unfiltered C4/noclean dataset. The plots show the average (and $95\%$ CI) of the aggregate performance metric on HellaSwag, CommonsenseQA, PIQA, SIQA and Wikipedia. The DataRater trained with eight $400$M inner models obtains the highest and most consistent performance across the two evaluated model scales and four meta-training checkpoints.
@@
We experimented with various hyper-parameter choices (including varying the meta-unroll length, the size of the population, etc.) and found that meta-evaluation performance depends mostly on the scale of the inner models in the population. In Figure 9 we show the relative performance improvement due to DataRater models trained on populations of inner models with different scales ($150$M and $400$M). We perform this experiment only on the C4/noclean underlying dataset and use a fixed discard fraction of $50\%$.
@@
In this plot (Fig. 9) we show relative performance improvement as a function of the meta-training checkpoint index (i.e., number of meta-update steps) and inner model scale. We calculate the relative performance improvement as the average percentage improvement in the raw metric (NLL for the C4/noclean validation set; accuracy for all the downstream tasks stated in the figure caption) of an LLM trained on the DataRater filtered data vs. a $1$B baseline LLM (trained like-for-like on the unfiltered C4/noclean dataset).
@@
We observe that the DataRater trained on a population of eight $400$M inner models results in the best performance across the board. So, we choose this configuration ($8$ x $400$M) and the meta-training checkpoint of $8000$ steps to produce all results in the main paper body – including for experiments which use other underlying training datasets (C4 and the Pile).
A.2 Experimental details.
@@
For evaluation, we train causal auto-regressive LLMs from random initialisation, following the Chincilla model architecture, on data filtered by a DataRater model. We use the Adam optimiser [Kingma and Ba, 2015] with a learning rate of $0.001$ for LLMs of size $50$M, $150$M and $400$M and one of $0.0002$ for $1$B models, with a standard cosine weight decay schedule. We train $50$M, $150$M, $400$M and $1$B models for $5$k, $12$k, $30$k and $48$k update steps respectively, with a $128$ batch size (except for $1$B for which we use a doubled batch size of $256$), with a sequence length of $2048$. Token budgets per model size can be inferred from the number of update steps, batch size and sequence length. For meta-training, for each inner LLM, we use a batch size of $128$ to compute the inner loss, and an outer batch size of $128$ to compute the outer loss. We used the Adam optimiser, with a decoupled weight decay of $0.1$ [Loshchilov and Hutter, 2019], with $100$ steps of linear warmup and global norm clipping [Pascanu et al., 2013] of $0.01$.
A.3 Infrastructure details.
@@
We implemented our infrastructure using the jax framework [Bradbury et al., 2018] and its ecosystem of associated libraries. Our meta-training framework is inspired by the Podracer architecture for distributed reinforcement learning [Hessel et al., 2021], and it enables parallel computation of meta-gradients from the inner LLM population. We have run our experiments on Google TPUs: we used $4\times 4\times 4$ topology v5 TPUs for meta-training, and up to $4\times 8$ topology v6e TPUs for evaluation. We carefully optimised our infrastructure, using data parallelism, model sharding as well as by using most techniques from MixFlow-MG [Kemaev et al., 2025] to keep RAM (i.e. HBM) consumption within hardware constraints.
A.4 Discussion on DataRater framework extensions.
@@ Appendix C Limitations
We demonstrated upward scale generalisation: from $400$M parameter inner models to $1$B models for evaluation (which are trained on DataRater-filtered data, not used as inner models themselves) – i.e., showing generalisation to LLMs which require an order of magnitude more FLOPS to train. As discussed above, we have also experimented with scaling up meta-gradient computation for larger models (i.e., using internal models with $27$B parameters) for the supervised fine-tuning regime, where inner model updates are made in a low-rank projection of parameter space using LORA [Hu et al., 2022]. This setting was non-trivial to tune for staying within hardware limitations on HBM consumption, but feasible. However, the scalability of meta-training DataRater models for extremely large foundation models with fully dense inner updates remains an open question, and may require further algorithmic advancements in scalable bilevel optimisation.

### v2

@@ 5 Conclusion
This approach offers a meaningful step towards automating and improving current dataset curation practices, which currently rely heavily on manual tuning and handcrafted heuristics. Our method provides a scalable and principled way to determine the value of individual data relative to the specified meta-objective, and shows considerable compute efficiency gains for filtering multiple real-world datasets when training on models of significant scale – thus providing a first proof of concept that meta-learning the data curation pipeline for modern LLMs is possible.
@@ Acknowledgements
The authors would like to express their gratitude to John E. Reid, Paul Michel, Katie Millican, Andrew Dai, Andrei A. Rusu, Jared Lichtarge, Chih-Kuan Yeh, Oriol Vinyals, the Gemini Data Team and the RL team for fruitful discussions, valuable feedback, and support throughout the project. The authors are also grateful to the anonymous reviewers for their insightful comments and constructive suggestions.
@@ Appendix A Appendix
Figure 9: Relative performance improvement induced by DataRater models trained on populations consisting of different sizes of inner models, as a function of the meta-training checkpoint. The underlying training dataset is C4/noclean and all evaluations are done with a $50\%$ discard fraction. The relative performance improvement is calculated with respect to a 1B baseline trained on the unfiltered C4/noclean dataset. The plots show the average (and $95\%$ CI) of the aggregate performance metric on HellaSwag, CommonsenseQA, PIQA, SIQA and Wikipedia. The DataRater trained with eight 400400M inner models obtains the highest and most consistent performance across the two evaluated model scales and four meta-training checkpoints.
@@
We experimented with various hyper-parameter choices (including varying the meta-unroll length, the size of the population, etc.) and found that meta-evaluation performance depends mostly on the scale of the inner models in the population. In Figure 9 we show the relative performance improvement due to DataRater models trained on populations of inner models with different scales (150150M and 400400M). We perform this experiment only on the C4/noclean underlying dataset and use a fixed discard fraction of $50\%$.
@@
In this plot (Fig. 9) we show relative performance improvement as a function of the meta-training checkpoint index (i.e., number of meta-update steps) and inner model scale. We calculate the relative performance improvement as the average percentage improvement in the raw metric (NLL for the C4/noclean validation set; accuracy for all the downstream tasks stated in the figure caption) of an LLM trained on the DataRater filtered data vs. a 11B baseline LLM (trained like-for-like on the unfiltered C4/noclean dataset).
@@
We observe that the DataRater trained on a population of eight 400400M inner models results in the best performance across the board. So, we choose this configuration ($8$ x 400400M) and the meta-training checkpoint of $8000$ steps to produce all results in the main paper body – including for experiments which use other underlying training datasets (C4 and the Pile).
A.2 Perplexity-based filtering.
We compare the DataRater to perplexity-based filtering, as perplexity is a common heuristic for assessing data quality that does not require choosing a specific validation set. For a fair and balanced comparison, all experiments in this section are conducted at the 150150M model scale. We use a 5050M DataRater model that was meta-trained on inner models of 150150M scale (as opposed to the 400400M inner models used for the results presented in the main paper).
We implement a perplexity-based filtering method similar to the one by Ankner et al. [2025]. To mirror the DataRater’s online, batch-level filtering mechanism, we first train three separate 150150M baseline models on held-out subsets of each of the three training sets used throughout this paper. Each of these baseline models then acts as a scorer, assigning a perplexity score to each data point (i.e. packed tokenized text sequence). Filtering is then performed within each oversampled batch by selecting data based on these scores. We use the same oversampling mechanism for filtering as for the DataRater method. We test the three variants of perplexity-based filtering introduced in [Ankner et al., 2025]: keeping only the data with the lowest perplexity (PPLX (Bot)), the highest perplexity (PPLX (Top)), and the data from the middle of the perplexity distribution within each batch (PPLX (Mid)).
Figure 10: Comparison of data filtering methods by sweeping over the filtering fraction. The plots show validation loss as a function of the proportion of data filtered out for the DataRater (150150M), there perplexity-based filtering baselines (PPLX-(Top, Mid, Bot)), and a no-filtering baselines. The PPLX (Mid) variant is the most effective of the perplexity-based methods, but the DataRater (150150M) achieves the best performance across the majority of filtering fractions.
Figure 11: Detailed comparison of DataRater (150150M) to the best perplexity-based filter PPLX (Mid). Baseline (no-filtering) performance is shown using dotted horizontal grey lines. This comparison uses the best filtering fractions for each method as identified from the sweep in Figure 10. Performance is evaluated across seven metrics for each of the three datasets. The DataRater (150150M) results in the best performance on 16/21 evaluations, while PPLX (Mid) shows better performance on four.
Figure 10 presents a sweep over different filtering fractions for the perplexity-based methods alongside the DataRater (150150M) and the no-filtering baseline on the three datasets. From this experiment we observe that PPLX (Mid) is the most effective of the perplexity-based filtering variants. The best validation set performance is achieved by filtering data using the DataRater (150150M), across the majority of the filtering fractions evaluated. Interestingly, filtering by perplexity can result in improvements when filtering C4 with very high filtering proportions. Based on this experiment, we select the best variant of perplexity-based filter as the PPLX (Mid) for the next comparison – together with the best filtering fractions: $90\%$ for the PPLX (Mid) method on all three datasets, and $75\%$, $50\%$ and $25\%$ for the DataRater (150150M) for the Pile, C4/noclean and C4 datasets respectively. In Figure 11 we present an extensive comparison of the selected methods on the seven evaluation metrics used in the main paper. This comparison shows that filtering using the DataRater still results in the best performance across most evaluations (i.e. 16 out of 21 evaluations). Perplexity-based filtering (specifically, keeping the “middle” perplexity data from each oversampled batch) shows better performance on four evaluations (SIQA on the Pile, ARC Easy and CommonsenseQA on C4/noclean, and PIQA on C4), while best performance on SIQA results from not filtering C4.
A.3 Meta-training stability.
As mentioned in the main paper body, we stabilise meta-gradient computation by using a population of inner models, passing their meta-gradients through individual Adam optimisers (averaging the resulting meta-gradient updates) and periodically reinitialising inner models. Since our goal is to use a trained and frozen DataRater to filter datasets for future downstream runs we require the DataRater to generalise across all stages of training – rather than to overfit to a particular inner model or to a particular stage of training. Having a population of inner models, whose parameters are reset occasionally, helps maintain diversity over the stage of training and promotes this type of generalisation. Our stabilisation strategies result in stable and bounded meta-updates as shown in Figure 12; the figure also highlights how inner models’ lifetimes are stratified. Furthermore, Figure 13 shows the correlation between DataRater scores across meta-training, demonstrating that meta-training starts converging after a few thousand steps when scores start tending towards perfect correlation (>0.95>0.95).
Figure 12: Meta-gradient update norms through meta-training. The plots show the L2 norm of meta updates originating from each individual inner model over the meta-training step counts of the C4/noclean 150M DataRater. The individual plots highlight when inner model parameters are reset by vertical dotted lines and by plotting individual lifetimes in different colours.
Figure 13: Correlation of DataRater scores between different stages of meta-training. This heatmap displays the Spearman correlation between DR scores (of the Pile 150M DR) obtained at various meta update checkpoints throughout training. All correlations are calculated using a static set of $25600$ examples from the Pile. The plot demonstrates that the DR scores stabilize after a few thousand meta-training steps. While scores from early checkpoints show weaker correlation with later ones (as well as each other), scores from checkpoints after approximately 3000 steps become almost perfectly correlated (correlation >0.95>0.95), indicating convergence.
A.4 Relation to heuristic filters.
We analyse the relationship between the DataRater score and common heuristics used for natural language data filtering. We adopted the heuristics used for filtering the C4 dataset, sourced from the DataTrove library [^3] https://github.com/huggingface/datatrove/blob/main/src/datatrove/pipeline/filters/c4_filters.py, along with other common metrics (e.g., non-alphanumeric character fraction, packed sub-sequence count, and word-level type-token ratio). A complete list of these heuristics is provided in Table 1.
To perform this analysis, we sampled $25600$ packed sequences from the Pile, calculated the DataRater score (using the the Pile 150M DR model) for each, and extracted all corresponding heuristic features. We first computed the Pearson correlation between the DataRater score and all the heuristics which is shown in Figure 14. This analysis shows that the DataRater score is most strongly correlated with features related to document size, such as the number of packed sub-sequences, characters, words, and sentences. Conversely, the score is anti-correlated with measures of non-standard content, such as the fractions of non-alphanumeric characters and punctuation. Notably, the DataRater score has only a weak correlation ($0.22$) with the composite C4 filter, c4_passes_all, which aggregates all individual C4 checks.
A simple linear regression model predicting the DataRater score from these features achieves a high $R^{2}$ of $0.766$. However, as many heuristics are highly collinear (i.e., strongly correlated or anti-correlated), this model is unsuitable for identifying the most predictive features. To isolate a minimal set of predictors, we instead fit a Lasso ($L_{1}$) regression model. We applied strong regularisation ($\alpha=0.03$) and normalised all features (Z-scoring). We trained the Lasso model with $10$-fold cross-validation (repeated $5$ times), achieving an $R^{2}$ of $0.753$. The resulting coefficients, shown in Figure 15, reveal that the Lasso model set most feature weights to zero, identifying only a few as significant predictors. The score is most strongly and positively predicted by the number of packed sub-sequences (num_articles, +0.42+0.42). The presence of curly braces (c4_curly_brace, +0.27+0.27) and greater text length (text_length_cutoff, +0.19+0.19; word_count, +0.13+0.13) also yield moderate positive coefficients. The only strong negative predictor is a high fraction of non-alphanumeric characters (0.33-0.33). In summary, the Lasso model indicates the DataRater score is primarily a function of document count (i.e., packed articles) and length, penalised by non-alphanumeric content. The Lasso model fit indicates that DataRater score is notably insensitive to most other common linguistic heuristics or C4-style filters.
Figure 14: Pearson correlation between DataRater score and various text filtering heuristics. The score is most strongly correlated with document size metrics and anti-correlated with non-standard content, such as non-alphanumeric text and punctuation. Notably, the correlation with the composite C4 filter, c4_passes_all, is low ($0.20$).
Figure 15: Coefficients of a Lasso ($L_{1}$) regression model fit to predict the DataRater score from Z-scored heuristic features. The model uses strong regularisation ($\alpha=0.03$) and achieves an $R^{2}$ of $0.753$ (vs. $0.766$ of OLS) assigning most features a zero coefficient. The DR score is most strongly predicted by the number of packed sub-sequences, the presence of curly braces and features related to the document length. The only strong negative predictor is the proportion of non-alphanumeric content.
A.5 Experimental details.
@@
For evaluation, we train causal auto-regressive LLMs from random initialisation, following the Chincilla model architecture, on data filtered by a DataRater model. We use the Adam optimiser [Kingma and Ba, 2015] with a learning rate of $0.001$ for LLMs of size 5050M, 150150M and 400400M and one of $0.0002$ for 11B models, with a standard cosine learning rate decay schedule. We train 5050M, 150150M, 400400M and 11B models for 55k, 1212k, 3030k and 4848k update steps respectively, with a $128$ batch size (except for 11B for which we use a doubled batch size of $256$), with a sequence length of $2048$. Token budgets per model size can be inferred from the number of update steps, batch size and sequence length. For meta-training, for each inner LLM, we use a batch size of $128$ to compute the inner loss, and an outer batch size of $128$ to compute the outer loss. We used the Adam optimiser, with a decoupled weight decay of $0.1$ [Loshchilov and Hutter, 2019], with $100$ steps of linear warmup and global norm clipping [Pascanu et al., 2013] of $0.01$.
A.6 Infrastructure details.
@@
We implemented our infrastructure using the jax framework [Bradbury et al., 2018] and its ecosystem of associated libraries [DeepMind et al., 2020]. Our meta-training framework is inspired by the Podracer architecture for distributed reinforcement learning [Hessel et al., 2021], and it enables parallel computation of meta-gradients from the inner LLM population. We have run our experiments on Google TPUs: we used $4\times 4\times 4$ topology v5 TPUs for meta-training, and up to $4\times 8$ topology v6e TPUs for evaluation. We carefully optimised our infrastructure, using data parallelism, model sharding as well as by using most techniques from MixFlow-MG [Kemaev et al., 2025] to keep RAM (i.e. HBM) consumption within hardware constraints.
A.7 Discussion on DataRater framework extensions.
@@ Appendix C Limitations
We demonstrated upward scale generalisation: from 400400M parameter inner models to 11B models for evaluation (which are trained on DataRater-filtered data, not used as inner models themselves) – i.e., showing generalisation to LLMs which require an order of magnitude more FLOPS to train. As discussed above, we have also experimented with scaling up meta-gradient computation for larger models (i.e., using inner models with 2727B parameters) for the supervised fine-tuning regime, where inner model updates are made in a low-rank projection of parameter space using LORA [Hu et al., 2022]. This setting was non-trivial to tune for staying within hardware limitations on HBM consumption, but feasible. However, the scalability of meta-training DataRater models for extremely large foundation models with fully dense inner updates remains an open question, and may require further algorithmic advancements in scalable bilevel optimisation.
Table 1: Description of text filtering heuristics use for correlation analysis.
Heuristic Description
packing_density Ratio of valid tokens to total tokens.
num_articles Number of packed sub-sequences.
text_length_cutoff Number of characters.
word_count Total word count (whitespace-delimited).
sentence_count Count of terminal punctuation (regex [.!?]).
[…]
unique_characters Ratio of unique characters to total text length.
word_type_token_ratio Word-level type-token ratio, i.e. number of unique words to number of words.
[…]
uppercase_letters_fraction Fraction of uppercase letters (A-Z).
[…]
average_word_length Mean word length (total non-whitespace chars / word_count).
numeric_fraction Fraction of numeric digits (0-9).
5gram_repetition_fraction Fraction of duplicate (non-unique) word-level 5-grams.
c4_lorem_ipsum True if the text contains "lorem ipsum".
c4_curly_brace True if the text contains a curly brace ({).
c4_contains_javascript True if any line contains the keyword "javascript".
c4_has_long_word True if any space-delimited word exceeds a length of 1000.
c4_fraction_lines_fail_terminal_punct Fraction of lines not ending with standard terminal punctuation (.?!"’).
c4_fraction_lines_fail_min_words Fraction of lines containing fewer than 3 words.
c4_contains_policy_keyword True if the text contains policy keywords (e.g., "privacy policy", "terms of use").
c4_sentence_count Sentence count derived from NLTK’s Punkt sentence tokenizer, assuming English text.
passes_all_c4_filters Composite filter: applies document-level (lorem ipsum, {) and line-level (length, punctuation, keywords) filtering, then checks for a minimum of five sentences.